OpenStudy (anonymous):

Each equation below represents a hyperbola. Which hyperbola will have a horizontal asymptote at y =7?

OpenStudy (anonymous):

a) y=5 over x+7 -4 b) y=5 over x-4 +7 c) y=5 over x-7 +4 d) y=5 over x-4 -7

OpenStudy (anonymous):

Help me please. I would really appreciate it.